3. An electron traveling horizontally enters a region where a uniform electric field is directed upward. What is the direction of the force exerted on the electron once it has entered the field? A) to the left D) downward electrou B) to the right E) out of the page, toward the reader C) upward
Answer #1

Electric force F = Eq = E*(-e)=-eE

from newtons law

acceleration a = F/m = -eE/m

the acceleration is down wards

the force acts downward